OpenStudy (anonymous):

find three points that solve the equation then plot them on the graph. x-3y=-6

OpenStudy (anonymous):

first get Y by itself....

OpenStudy (anonymous):

X-3Y=6 X-6=3Y X/3-2=Y

OpenStudy (anonymous):

how i did that is the equation you gave me is X-3Y=6. well to get Y by its self the y need to look like this that the end: X-/+M=Y. the Y also has to be positive. in this case you see that the Y is -3y. that -3y has to become +3y. so to do that you do this:  when you move anything to the opposite side of the = the sign changes. so now the problem looks like this now: x-6=3y now you divide like this: 3y/3, x/3, -6/3 then it'll look like this: x/3-2=y graph that and just find three points